# Disentangling Boltzmann brains, the time-asymmetry of memory, the H theorem, and the second law

### v2

@@ III.4 The Past Hypothesis and the Second Law
To see this, first we introduce a fully formal definition of the PH as another logically independent set of 3 suppositions:
@@
    The H theorem holds for a kernel generating a Boltzmann process ${\cal{X}}$;
@@
Suppose we had data and adopted associated priors that led us to conclude that items (i) and (ii) of the PH hold. For example, this could be because we have data ${\cal{D}}_{1}$ and ${\cal{D}}_{3}(2)$, and assume that both are reliable – but did not have any data ${\cal{D}}_{3}(1)$. Suppose that in addition we simply assumed that item (iii) of the PH holds, by incorporating that assumption into our prior. (While cosmological arguments indicate that that supposition holds, there are subtleties that are irrelevant for current purposes.)
If we made no other assumptions and had no other data, we would conclude that the stochastic process of the values of entropy between the Big Bang and now was monotonically increasing in expectation value. (More precisely, that data together with those assumptions would mean that the expected value of the marginal distributions of that process at single times $t$ increased monotonically with $t\in[t^{*},t_{0}]$.) So a prior that ${\cal{D}}_{1}$ and ${\cal{D}}_{3}(2)$ are reliable, together with the assumption that item (iii) of the PH holds, would imply the second law.
Note that the second law would in turn mean that entropy decreases the further back we go into our past from the present. However, as mentioned above, the BB hypothesis would imply that entropy increases going backward in time from the present. So the PH would both imply the second law and rule out the BB hypothesis.
@@
Of course, it is possible to combine the suppositions of both hypotheses, simply by supposing the full data set ${\cal{D}}_{3}$, including both subsets, is reliable. Doing this would lead us to conclude that the evolution of the universe’s entropy is equivalent to a Boltzmann process nailed to the pair $\{(t^{*},S^{*}),(t_{0},S_{0})\}$, as illustrated in Fig. 1. Reference Scharnhorst et al. (2024) analyzes such Boltzmann processes that are nailed to such a pair of time-entropy pairs. It is shown there that for certain parameters of the Boltzmann process and certain values in the nailed data set, the associated posterior expected entropy would have been increasing as we start to go backwards in time, into our very recent past. This would imply that the second law is wrong in our recent past. Expected entropy would still decrease though once we go far enough backward in time towards the Big Bang. So the second law would be correct on long enough timescales into our past. Moreover, we would still have the usual second law going forward from the present. So all experimental tests of the second law we might do in our future would (according to the H theorem, etc.) still confirm the second law. Such a scenario could be viewed as a highly attenuated variant of the BB hypothesis, one not necessarily having any implications about carefully arranged sensory inputs to our brains and / or carefully arranged photons streaming in to our telescopes, etc.
III.5 What the H theorem does not imply
@@
So the second law is consistent with the priors that imply it — but only by virtue of the fact that deriving it relies on circular reasoning. (See Section B.1 for a Bayesian formalization of circular reasoning, and discussion from a Bayesian perspective of the attendant flaws.) We can adopt priors that ${\cal{D}}_{1}$ and ${\cal{D}}_{3}$ are reliable, or that the second law holds (or even both). But we cannot somehow avoid making any prior assumptions at all, in an attempt to bootstrap ourselves into placing high probability on having ${\cal{D}}_{1}$ or ${\cal{D}}_{3}$ be reliable and that the second law of thermodynamics holds.
Suppose that we misunderstood the foregoing, and thought that somehow the reliability of ${\cal{D}}_{1}$ not only implies the H theorem, but also implies:
@@
Suppose as well that the second law is true. That supposition would mean (by (b)) that ${\cal{D}}_{1}$ is reliable, and therefore Newton’s laws hold. This in turn would mean (as Boltzmann showed) that the H theorem holds. By (a), we could then conclude that the BB hypothesis holds.
However, as pointed out above, this derivation of the BB hypothesis would mean that the second law does not hold. That would contradict the second supposition (b) of this very derivation of the BB hypothesis. Note that this is more than using circular reasoning to try to establish some hypothesis or other. Rather it is using reasoning that contradicts itself to try to establish a particular hypothesis. It is this seeming self-contradiction that people have had in mind when they have criticized the arguments for the BB hypothesis as being “cognitively unstable” Carroll (2020).
Of course, the foregoing does not mean that the BB hypothesis is false — it just means that the reasoning for that hypothesis given above contradicts itself. As elaborated above, reasoning based on the correct understanding of the H theorem simply has nothing to say about the BB hypothesis. This is our fourth contribution.
Even if this derivation of the BB hypothesis is cognitively unstable, that provides no reason to conclude that the probability of being a BB is low. Indeed, it is straightforward to construct a simple variant of the BB hypothesis that does not suffer from any form of “cognitive instability”.
This variant of the BB hypothesis is just like the standard BB hypothesis in that it supposes that the universe fluctuates from a state of maximal entropy down to some minimal entropy value, after which the entropy of the universe starts expanding again. The difference is that rather than suppose that the minimum of the entropy of the universe occurs at the present, under the 1000CE BB hypothesis it occurs at 1000CE. [^5] Such a scenario has sometimes been considered informally in the literature, where it has been called a “Boltzmann bubble” Chen (2022); Wallace (2023b). More formally, the “1000CE BB hypothesis” is the same three suppositions that were used to define the BB hypothesis in Section III.3, except that $t_{0}$ is changed to be the year 1000CE:
@@
    The H theorem holds for a kernel generating a Boltzmann process ${\cal{X}}$;
@@
Under the 1000CE BB hypothesis, the second law would hold, so our memory systems (scientific records) would accurately record the results of all experiments conducted in the last thousand years. So the data set ${\cal{D}}_{1}$ which implies the H theorem would be reliable. So cognitive instability arguments don’t apply, even though a version of the BB hypothesis is still true. [^6] There are other problems with the criticism that the justification of the BB hypothesis based on (a), (b) is unstable. Most obviously, these criticisms make their case by first seemingly proving some particular proposition $A$ based on some particular data (in out case, proving the H theorem from ${\cal{D}}_{1}$ via Newton’s laws). They then seemingly show that $A\rightarrow B$ (in our case $B$ is the BB hypothesis). They end by seemingly showing that $B$ contradicts $A$. From that seeming contradiction they conclude that $B$ must be wrong. That conclusion is fallacious; unless one can show explicitly that $A\rightarrow B$ is not logically sound, then the contradiction means that the premise $A$ must be wrong, with no implications for whether $B$ is wrong. In our case, this would mean that the contradiction proves that Newton’s laws are wrong. (!)
@@
Note that the 1000CE BB hypothesis would appear to be data-independent — no data could either confirm or refute it. And as we described in the introduction, there are an infinite number of hypotheses that appear to have this data-independent character, corresponding to different times of the entropy minimum. This is related to our comment in the introduction that there is a “grey mass” of hypotheses that appear to be outside the purview of science. However, note that we could keep pushing the time of the entropy minimum further and further back into the past, from the present to 1000CE and ultimately all the back to the big bang. At that point we have a variant of the BB hypothesis which is just the fluctuation hypothesis version of the PH Chen (2022). However, as discussed above, we do have a data set – ${\cal{D}}_{3}$ – that we think gives us some confirmation that entropy was indeed quite low at the big bang. So one must be careful before making sweeping claims about whether any particular variations of the BB hypothesis could or could not be confirmed by data. These formal definitions of the BB hypothesis, the PH, and the 1000CE hypothesis are our last contribution.
@@
As a final comment, as pointed out above, this same kind of data problem holds for the second law, the PH, etc., in that none of them can be either confirmed or refuted solely from data. Any data-based assessment of any of them ultimately relies on making assumptions for prior probabilities. In our analysis in this paper, we have chosen to only adopt the prior that ${\cal{D}}_{1}$ is reliable, which we take to be a necessary minimal assumption for any investigation of these issues. [^7] In particular, we have been careful not to introduce any assumptions about what the nailed set should be when we use the H theorem. The second law, the validity of scientific records, the PH, the BB hypothesis (in all of its invariants) all make some assumption for that nailed data set, without any formal justification in terms of Bayesian decision theory. Of course, this prior removes the need to invoke the second law to establish the reliability of ${\cal{D}}_{1}$. So there is no cognitive instability under this prior.
@@ IV Discussion
In this paper, we introduced new stochastic process theory to allow us to disentangle the second law, the reliability of human memory, the BB hypothesis, and the Past Hypothesis (PH) with full formal rigor.
@@
    Boltzmann’s H theorem, properly formalized, says that the H function evolves as a particular type of stochastic process, a “Boltzmann process”, and then derive some new, useful properties of Boltzmann processes;
@@
    We point out that due to (iii), the H theorem has nothing to say one way or another about the legitimacy of the BB hypothesis, nor of the second law, the PH, or the 1000CE hypothesis.
We then used these preliminary results to disentangle the BB hypothesis, the time-asymmetry of memory, the H theorem, and the second law, with more formal rigor than has been done before. Specifically, we demonstrate that neither the reliability of experimental data nor our possessing Type-3 memory can be disentangled from the second law. Second, we show that the validity of the H theorem is independent of the BB hypothesis, the PH, and the 1000CE hypothesis, since by it is an independent ingredient in those hypotheses. Third, we show that those hypotheses are all variants of one another, simply differing in the times they assume for the extreme of an entropy fluctuation — with the caveat that it the BB hypothesis has a seeming self-contradiction baked in.
@@
We will emphasize that since the H theorem cannot make a conclusion regarding probability of the BB hypothesis that doesn’t resort to priors (see Appendix), any conclusion that the BB hypothesis should be discarded entirely must move to the level of theory building and/or justification of priors. And as stated previously, it seems very hard to justify a prior that allows one to conclude that the BB hypothesis is true. In the context of theory building, philosophy of science has provided no shortage of qualities that might help define a “good” theory. Cognitive stability Carroll (2020) is one such quality. Yet, critically, there is not a fully rigorous argument, yet, in the domain of physics to dispel the possibility of the BB hypothesis. The view we have established in this paper is that the BB hypothesis is not established by the standard argument, nor is it refuted by standard criticisms. It seems the BB hypothesis falls out of the domain of science, into the grey ocean of hypothesis described in the introduction. We cannot rigorously assign probability to something unknowable, nor can we say that something unknowable is not true.
It is important to appreciate that there are many deep issues in the philosophy of science that our careful investigation of the BB hypothesis uncovers. One is that there is the huge issue of how — from the perspectives of either philosophy or the axiomatizations of Bayesian probability — one can justify conditioning on any moment in addition to the present, and if so, why just one such additional moment. Suppose that we do though, e.g., due to associated assumptions concerning reliability. This would still leave the issue of whether (as implied by the construction rule) we really should nail stochastic processes to all time-entropy pairs that we have very strong reason to believe. All of which in fact provides ways to rescue the BB hypothesis, by justifying a prior that we nail the Boltzmann process to only $(t_{0},S_{0})$.
@@ Acknowledgements.
Great thanks to Sean Carroll and Wayne Myrvold for many conversations on this topic. DHW also thanks the SFI for support.
@@ B.2 Unstable reasoning in analysis of Boltzmann Brains
where $B$ and $A$ are two random variables. Such a pair of conditional distributions is summarized as “unstable reasoning”.
@@
Combining that assumption with our conclusion that $P(A=a)\ll P(B=b)$ would mean that $P(B=b)\gg 1$, i.e., it would result in a contradiction. So there is no joint distribution $P(A,B)$ that obeys Eqs. 15, 16 and 19. Our supposition for at least one of those three distributions must be wrong.
@@
Eq. 15 would instead start from the fact that if ${\cal{D}}_{1}$ is reliable, then the second law holds, and so the H theorem does, i.e., the universe evolves as a Boltzmann process. Next, if we apply the norms of Bayesian reasoning, we must condition that stochastic process on the data we directly have, and nothing else. In other words, we should fix the Boltzmann process to the only moment about which we directly have data, namely the present. So we recover the BB hypothesis. In fact, this is precisely the standard argument for why the BB hypothesis must be true.
@@
In the literature, the BB hypothesis is commonly formulated this way, and sometimes combined with Eqs. 15 and 16. The resultant reasoning is sometimes referred to as “unstable”.
@@
If one assumed nothing more than Eqs. 15 and 16 though, we could only conclude that the probability of the BB hypothesis being true is far greater than the probability that ${\cal{D}}_{1}$ is reliable. In particular we could not conclude the value of $B$ is very likely to be the bit, “not true”. In other words, this pair of assumptions does not mean the BB hypothesis is likely to be false.
@@
Commonly in the literature people want to go further than just those two assumptions, and assume that ${\cal{D}}_{1}$ is highly likely to be reliable. Doing this though just means we have an instance of Eqs. 15, 16 and 19. Again, we cannot conclude that the value of $B$ must be the bit, “not true”. Instead we can only conclude that in point of fact, at least one of our three suppositions for the probabilities must be wrong.
@@
In light of these nuances, the definition of the BB hypothesis in the main text is far more careful. It does not presume that Eqs. 15 and 16 hold and / or that Eq. 19 also holds.
@@
As mentioned above, it is quite challenging to tease out the physics that determines which of those three suppositions is wrong. We can illustrate this issue informally, focusing on the assumption that $P(A=a\,|\,B=b)\simeq 0$. First note that the H theorem relies on some very strong assumptions. One of the most glaring is that it assumes that the system in question is closed to the outside world during its evolution. This assumption might hold for the case where the system in question is the entire universe (though even that assumption is quite problematic). But it certainly does not hold for any memory system recording the results of experiments we once performed — by definition, such a memory system is open to the outside world.
@@
The entropy of such a system can evolve in the opposite direction from that of the universe as a whole, the H theorem notwithstanding. [^8] Indeed, the standard thermodynamic understanding of modern day biological organisms is that they are subsystems of the universe which are tightly coupled to the outside world, in a special way that causes their entropy to evolve in the opposite direction of the universe as a whole. Under the BB hypothesis, subsystems whose entropy evolves in the opposite direction of that of the universe as a whole are equally likely to occur in our past as in our future. So we could have the entropy of the universe as a whole increase as one moves further into the past from our present, while the entropy of a subsystem of the universe which is coupled to the outside world decreases as one moves further into the past from our present. Given this, suppose the BB hypothesis holds, and consider our memory system whose current contents are ${\cal{D}}_{1}$, which was coupled to a set of experimental tests of Newton’s laws sometime in our past. Since it has been tightly coupled with the outside world, the entropy of that memory system might be decreasing into our past even as the entropy of the entire universe increases into our past. In turn, the decrease of entropy of the memory system into our past is precisely what one needs for the contents of that memory to be reliable, according to the standard thermodynamic processes that establish the reliability of memory Wolpert and Kipper (2024); Rovelli (2022). In such a scenario, it would not be the case that $P(A=a\,|\,B=b)\simeq 0$.

### v3

@@ III.4 The Past Hypothesis and the Second Law
To see this, first we introduce a fully formal definition of the PH, just like our formal definition of the BB hypothesis, as a set of 3 logically independent suppositions:
@@
    The entropy conjecture holds for a kernel generating a Boltzmann process $\cal{X}$};
@@
Note that this formal definition of the PH is almost identical to the formal definition of the BB hypothesis, with only the second of the three assumptions being (slightly) different in the two cases.
To begin, in direct analogy to the case with the BB hypothesis, simply assume that item (iii) of the PH holds, by incorporating that assumption into our prior. Next, suppose we had data and adopted associated priors that led us to conclude that item (i) of the PH holds, and that the nailed data set should include $(t^{*},S^{*})$. For example, this could be because we have data ${\cal{D}}_{1}$ and ${\cal{D}}_{3}(2)$, and assume that both are reliable. Suppose though that we did not have any data ${\cal{D}}_{3}(1)$.
If we made no other assumptions and had no other data, then by the construction rule, we would conclude that the stochastic process of the values of entropy was a Boltzmann process singleton-equivalent to $(t^{*},S^{*})$. In other words, it would establish item (ii) of the PH, to go with items (i) and (iii). So the PH would hold in its entirety.
That in turn would lead us to conclude that the values of entropy between the Big Bang and now was monotonically increasing in expectation value, and that it would continue to increase into our future. (More precisely, that data together with those assumptions would mean that the expected value of the marginal distributions of that process at single times $t$ increased monotonically with $t\in[t^{*},t_{0}]$.)
So a prior that ${\cal{D}}_{1}$ and ${\cal{D}}_{3}(2)$ are reliable, together with the assumption that item (iii) of the PH holds, and no other data, would imply that the PH holds, and that therefore the second law holds. Note that the second law would in turn mean that entropy decreases the further back we go into our past from the present. However, as mentioned above, the BB hypothesis would imply that entropy increases going backward in time from the present. So the PH would both imply the second law and rule out the BB hypothesis.
@@
Of course, it is possible to combine the suppositions of both hypotheses, simply by supposing the full data set ${\cal{D}}_{3}$, including both subsets, is reliable. Doing this would lead us to conclude that the evolution of the universe’s entropy is equivalent to a Boltzmann process nailed to the pair $\{(t^{*},S^{*}),(t_{0},S_{0})\}$, as illustrated in Fig. 1. Reference Scharnhorst et al. (2024) analyzes such Boltzmann processes that are nailed to such a pair of time-entropy pairs. It is shown there that for certain parameters of the Boltzmann process and certain values in the nailed data set, the associated posterior expected entropy would have been increasing as we start to go backwards in time, into our very recent past. This would imply that the second law is wrong in our recent past. Expected entropy would still decrease though once we go far enough backward in time towards the Big Bang. So the second law would be correct on long enough timescales into our past. Moreover, we would still have the usual second law going forward from the present. So all experimental tests of the second law we might do in our future would (according to the entropy conjecture, etc.) still confirm the second law. Such a scenario could be viewed as a highly attenuated variant of the BB hypothesis, one not necessarily having any implications about carefully arranged sensory inputs to our brains and / or carefully arranged photons streaming in to our telescopes, etc.
III.5 What the entropy conjecture does not imply
@@
So the second law is consistent with the priors that imply it — but only by virtue of the fact that deriving it relies on circular reasoning. We can adopt priors that ${\cal{D}}_{1}$ and ${\cal{D}}_{3}$ are reliable, or that the second law holds (or even both). But we cannot somehow avoid making any prior assumptions at all, in an attempt to bootstrap ourselves into placing high probability on having ${\cal{D}}_{1}$ or ${\cal{D}}_{3}$ be reliable and that the second law of thermodynamics holds. (See Section B.1 for a Bayesian formalization of circular reasoning, and associated derivation of what circular reasoning does and does not imply.)
Suppose that we misunderstood the foregoing, and thought that somehow the reliability of ${\cal{D}}_{1}$ not only implies the entropy conjecture, but also implies:
@@
Suppose as well that the second law is true. That supposition would mean (by (b)) that ${\cal{D}}_{1}$ is reliable, and therefore Newton’s laws hold. This in turn would mean (as Boltzmann showed) that the entropy conjecture holds. By (a), we could then conclude that the BB hypothesis holds.
However, this derivation of the BB hypothesis would mean that the second law does not hold. That would contradict the second supposition (b) of this very derivation of the BB hypothesis. Note that this is more than using circular reasoning to try to establish some hypothesis or other. Rather it is using reasoning that contradicts itself to try to establish a particular hypothesis. It is this seeming self-contradiction that we presented in Section II, and is related to unstable reasoning.
Of course, the foregoing flaw in a particular derivation of the BB hypothesis does not mean that the BB hypothesis is false — it just means that the reasoning for that hypothesis given above contradicts itself. More precisely, as we prove in Section B.1, unstable reasoning just means that one is implicitly trying to use a joint prior distribution over all associated random variables that is not a proper probability distribution. So it’s simply saying that one needs to adopt a different prior. This clarification that issues of unstable reasoning have no bearing on the BB hypothesis is our fifth contribution.
[…]
It is illuminating to push this scenario further, by introducing a slight variant of the BB hypothesis. This variant is just like the standard BB hypothesis in that it supposes that the universe fluctuates from a state of maximal entropy down to some minimal entropy value, after which the entropy of the universe starts expanding again. The difference is that rather than suppose that the minimum of the entropy of the universe occurs at the present, under the 1000CE BB hypothesis it occurs at 1000CE. [^6] Such a scenario has sometimes been considered informally in the literature, where it has been called a “Boltzmann bubble” Chen (2022); Wallace (2023b). More formally, the “1000CE BB hypothesis” is the same three suppositions that were used to define the BB hypothesis in Section III.3, except that $t_{0}$ is changed to be the year 1000CE:
@@
    The entropy conjecture holds for a kernel generating a Boltzmann process $\cal{X}$};
@@
Under the 1000CE BB hypothesis, the second law would hold, so our memory systems (scientific records) would accurately record the results of all experiments conducted in the last thousand years. So the data set ${\cal{D}}_{1}$ which implies the entropy conjecture would be reliable. So cognitive instability arguments don’t apply, even though a version of the BB hypothesis is still true. [^7] There are other problems with the criticism that the justification of the BB hypothesis based on (a), (b) is unstable. Most obviously, these criticisms make their case by first seemingly proving some particular proposition $A$ based on some particular data (in out case, proving the entropy conjecture from ${\cal{D}}_{1}$ via Newton’s laws). They then seemingly show that $A\rightarrow B$ (in our case $B$ is the BB hypothesis). They end by seemingly showing that $B$ contradicts $A$. From that seeming contradiction they conclude that $B$ must be wrong. That conclusion is fallacious; unless one can show explicitly that $A\rightarrow B$ is not logically sound, then the contradiction means that the premise $A$ must be wrong, with no implications for whether $B$ is wrong. In our case, this would mean that the contradiction proves that Newton’s laws are wrong. (!)
@@
Note that we could keep pushing the time of the entropy minimum further and further back into the past, from the present to 1000CE and ultimately all the back to the big bang. At that point we have a variant of the BB hypothesis which is just the fluctuation hypothesis version of the PH Chen (2022). Moreover, as discussed above, we do have a data set – ${\cal{D}}_{3}$ – that we think gives us some confirmation that entropy was indeed quite low at the big bang. So one must be careful before making sweeping claims about whether any particular variations of the BB hypothesis could or could not be confirmed by data, since the PH is itself only such a variant. These formal definitions of the BB hypothesis, the PH, and the 1000CE hypothesis are our last contribution.
@@
As a final comment, as pointed out above, this same kind of data problem holds for the second law, the PH, etc., in that none of them can be either confirmed or refuted solely from data. Any data-based assessment of any of them ultimately relies on making assumptions for prior probabilities. In our analysis in this paper, we have chosen to only adopt the prior that ${\cal{D}}_{1}$ is reliable, which we take to be a necessary minimal assumption for any investigation of these issues. [^8] In particular, we have been careful not to introduce any assumptions about what the nailed set should be when we use the entropy conjecture. The second law, the validity of scientific records, the PH, the BB hypothesis (in all of its invariants) all make some assumption for that nailed data set, without any formal justification in terms of Bayesian decision theory. Of course, this prior removes the need to invoke the second law to establish the reliability of ${\cal{D}}_{1}$. So there is no cognitive instability under this prior.
@@ IV Discussion
In this paper, we extended recent advances in stochastic process theory in order to disentangle the second law, the reliability of human memory, the BB hypothesis, and the Past Hypothesis (PH) with full formal rigor.
@@
    Boltzmann’s entropy conjecture, properly formalized, says that the H function evolves as a particular type of stochastic process, a “Boltzmann process”, and then derive some new, useful properties of Boltzmann processes;
@@
    We point out that due to (iii), the entropy conjecture has nothing to say one way or another about the legitimacy of the BB hypothesis, nor of the second law, the PH, or the 1000CE hypothesis.
We then used these preliminary results to disentangle the BB hypothesis, the time-asymmetry of memory, the entropy conjecture, and the second law, with more formal rigor than has been done before. Specifically, we demonstrate that neither the reliability of experimental data nor our possessing Type-3 memory can be disentangled from the second law. Second, we show that the validity of the entropy conjecture is independent of the BB hypothesis, the PH, and the 1000CE hypothesis, since it is an independent ingredient in those hypotheses. Third, we show that those hypotheses are all variants of one another, simply differing in the times they assume for the extreme of an entropy fluctuation — with the caveat that it the BB hypothesis has a seeming self-contradiction baked in.
@@
We will emphasize that since the entropy conjecture cannot make a conclusion regarding probability of the BB hypothesis that doesn’t resort to priors (see Appendix), any conclusion that the BB hypothesis should be discarded entirely must move to the level of theory building and/or justification of priors. And as stated previously, it seems very hard to justify a prior that allows one to conclude that the BB hypothesis is true. In the context of theory building, philosophy of science has provided no shortage of qualities that might help define a “good” theory. Cognitive stability Carroll (2020) is one such quality. Yet, critically, there is not a fully rigorous argument, yet, in the domain of physics to dispel the possibility of the BB hypothesis. The view we have established in this paper is that the BB hypothesis is not established by the standard argument, nor is it refuted by standard criticisms. It seems the BB hypothesis falls out of the domain of science, into the grey ocean of hypothesis described in the introduction. We cannot rigorously assign probability to something unknowable, nor can we say that something unknowable is not true.
It is important to appreciate that there are many deep issues in the philosophy of science that our careful investigation of the BB hypothesis uncovers. One is that there is the huge issue of how — from the perspectives of either philosophy or the axiomatizations of Bayesian probability — one can justify conditioning on any moment in addition to the present, and if so, why just one such additional moment.
Suppose that we do though, e.g., due to associated assumptions concerning reliability. This would still leave the issue of whether (as implied by the construction rule) we really should nail stochastic processes to all time-entropy pairs that we have very strong reason to believe. All of which in fact provides ways to rescue the BB hypothesis, by justifying a prior that we nail the Boltzmann process to only $(t_{0},S_{0})$.
@@ Acknowledgements.
Great thanks to Sean Carroll and Wayne Myrvold for many conversations on this topic, and thanks also to Artemy Kolchinsky. DHW also thanks the SFI for support.
@@ B.2 Unstable reasoning in analysis of Boltzmann Brains
where BB and $A$ are two random variables. Such a pair of conditional distributions is summarized as “unstable reasoning”.
@@
Combining that assumption with our conclusion that $P(A=a)\ll P(B=b)$ would mean that $P(B=b)\gg 1$, i.e., it would result in a contradiction. So there is no joint distribution $P(A,B)$ that obeys Eqs. 12, 13 and 16. Our supposition for at least one of those three distributions must be wrong.
@@
Eq. 12 would instead start from the fact that if ${\cal{D}}_{1}$ is reliable, then the second law holds, and so the entropy conjecture does, i.e., the universe evolves as a Boltzmann process. Next, if we apply the norms of Bayesian reasoning, we must condition that stochastic process on the data we directly have, and nothing else. In other words, we should fix the Boltzmann process to the only moment about which we directly have data, namely the present. So we recover the BB hypothesis. In fact, this is precisely the standard argument for why the BB hypothesis must be true.
@@
In the literature, the BB hypothesis is commonly formulated this way, and sometimes combined with Eqs. 12 and 13. The resultant reasoning is sometimes referred to as “unstable”.
@@
If one assumed nothing more than Eqs. 12 and 13 though, we could only conclude that the probability of the BB hypothesis being true is far greater than the probability that ${\cal{D}}_{1}$ is reliable. In particular we could not conclude the value of $B$ is very likely to be the bit, “not true”. In other words, this pair of assumptions does not mean the BB hypothesis is likely to be false.
@@
Commonly in the literature people want to go further than just those two assumptions, and assume that ${\cal{D}}_{1}$ is highly likely to be reliable. Doing this though just means we have an instance of Eqs. 12, 13 and 16. Again, we cannot conclude that the value of $B$ must be the bit, “not true”. Instead we can only conclude that in point of fact, at least one of our three suppositions for the probabilities must be wrong.
@@
In light of these nuances, the definition of the BB hypothesis in the main text is far more careful. It does not presume that Eqs. 12 and 13 hold and / or that Eq. 16 also holds.
@@
As mentioned above, it is quite challenging to tease out the physics that determines which of those three suppositions is wrong. We can illustrate this issue informally, focusing on the assumption that $P(A=a\,|\,B=b)\simeq 0$. First note that the entropy conjecture relies on some very strong assumptions. One of the most glaring is that it assumes that the system in question is closed to the outside world during its evolution. This assumption might hold for the case where the system in question is the entire universe (though even that assumption is quite problematic). But it certainly does not hold for any memory system recording the results of experiments we once performed — by definition, such a memory system is open to the outside world.
@@
The entropy of such a system can evolve in the opposite direction from that of the universe as a whole, the entropy conjecture notwithstanding. [^9] Indeed, the standard thermodynamic understanding of modern day biological organisms is that they are subsystems of the universe which are tightly coupled to the outside world, in a special way that causes their entropy to evolve in the opposite direction of the universe as a whole. Under the BB hypothesis, subsystems whose entropy evolves in the opposite direction of that of the universe as a whole are equally likely to occur in our past as in our future. So we could have the entropy of the universe as a whole increase as one moves further into the past from our present, while the entropy of a subsystem of the universe which is coupled to the outside world decreases as one moves further into the past from our present. Given this, suppose the BB hypothesis holds, and consider our memory system whose current contents are ${\cal{D}}_{1}$, which was coupled to a set of experimental tests of Newton’s laws sometime in our past. Since it has been tightly coupled with the outside world, the entropy of that memory system might be decreasing into our past even as the entropy of the entire universe increases into our past. In turn, the decrease of entropy of the memory system into our past is precisely what one needs for the contents of that memory to be reliable, according to the standard thermodynamic processes that establish the reliability of memory Wolpert and Kipper (2024); Rovelli (2022). In such a scenario, it would not be the case that $P(A=a\,|\,B=b)\simeq 0$.